# Effective Whitney Stratification of Real Algebraic Varieties

### v3

@@ 14 Return $(X_{\bullet},\mathbf{S}_{\bullet}X)$
The termination of WTBStratFlag follows from noting that the recursive call in Line 9 involves the variety $X_{d}$ of dimension at most $d$, with $d$ strictly smaller than $k=\dim X$. And the fact that $X_{\bullet}$ is a Whitney stratification of $X$ follows from the correctness of the original WhitStrat algorithm [26, Sec 4.4]. Moreover, since $\mathbf{F}_{\bullet}X$ is merged into $\mathbf{S}_{\bullet}X$ via Line 7, verifying that $X_{\bullet}$ is $\mathbf{F}_{\bullet}$-subordinate can be reduced to checking that it is $\mathbf{S}_{\bullet}$-subordinate. Therefore, it suffices to confirm that the output pair $(X_{\bullet},\mathbf{S}_{\bullet}X)$ satisfies the two requirements of Definition 2.6. To this end, consider an $i$-dimensional stratum $M$ of $X_{\bullet}$.
@@ 8         Set $\mathbf{F}_{i}X:=f^{-1}(Y^{\prime}_{i})$
10 Set $(X_{\bullet},Y_{\bullet}):=$ Refine$(X^{\prime}_{\bullet},Y^{\prime}_{\bullet},f)$
@@ 11 Return $(X_{\bullet},Y_{\bullet})$
Noting that of $X^{\prime}_{\bullet}$ is subordinate to the flag $\mathbf{F}_{\bullet}$ by Line 9 of WhitStratMap, we know that for each stratum $M$ of $X^{\prime}_{\bullet}$ there is a number $i:=i(M)$ satisfying $M\subset(\mathbf{F}_{i}X-\mathbf{F}_{i-1}X)$. Now by Line 8, for any such stratum we have $f(M)\subset(Y^{\prime}_{i}-Y^{\prime}_{i-1})$. By the definition of a Whitney stratum we know that $M$ is connected, and so its image under the continuous map $f$ must also be connected; thus there is a unique stratum $N\subset(Y^{\prime}_{i}-Y^{\prime}_{i-1})$ of $Y^{\prime}_{\bullet}$ satisfying $f(M)\subset N$, as claimed above. It remains to establish that the Jacobian of $f|_{M}:M\to N$ has maximal rank at every point of $x$.
@@
Recall that $\mathbf{S}^{\prime\prime}_{j}X$ contains $\mathbf{R}_{j}\overline{M}$ by the second requirement of 2.6 (and similarly for $j-1$). Therefore, the restricted map $f^{*}:\overline{M}_{\rm reg}\to\mathbb{C}^{m}$ has Jacobian of constant rank $\rho_{j}$ given by
@@
Since $M$ is an open subset of $\overline{M}_{\rm reg}$, for each $x\in M$ we immediately have rank ($\text{rank }\big{(}\text{\bf J}f^{*}(x)\big{)}=\rho_{j}.$ Let $\iota:N\hookrightarrow\mathbb{C}^{m}$ be the inclusion map, and recall that $f|_{M}:M\to N$ denotes the restriction of $f$ to $M$, now viewed as a smooth map to $N$ rather than to $\mathbb{C}^{m}$. Thus, $f^{*}=\iota\circ f|_{M}$, and the following diagram of derivatives commutes by the chain rule:
@@ 4         Add $\overline{f(S)}$ to $Y^{+}_{\geq d}$
 5         Merge $Y^{+}_{\bullet}$ with WhitStrat$(\text{\rm Pure}_{d}(Y^{\prime}_{d}))$
@@ 10                           Add $Z$ to $X^{\prime}_{\geq r}$
11                           Merge $X^{\prime}_{\bullet}$ with WTBStrat$({\rm Pure}_{r}(X^{\prime}_{r}),f)$
@@ 14 Return $(X^{\prime}_{\bullet},Y^{\prime}_{\bullet})$
From Theorem 2.7 and Proposition 2.8, we obtain Proposition 2.9 below, which is analogous to [26, Proposition 6.7]. The novelty here lies in the fact that the result below applies to arbitrary polynomial maps between affine complex varieties, and does not impose any genericity requirements. The proof, however, is essentially identical to that of [26, Proposition 6.7].
@@
In [26, Proposition 6.7] this result is obtained for generic mappings $f:X\to Y$, as defined in [26, Definition 6.3], using the construction of the Refine subroutine along with a result ([26, Proposition 6.6]) analogous to that of Proposition 2.8 in the case of generic maps. The new Refine subroutine above is identical to that in [26, Section 6] other than the new version uses WTBStrat at Line 11, rather than WhitStrat. The argument is then identical to that of [26, Proposition 6.7], except we employ the more general result of Proposition 2.8 in place of [26, Proposition 6.6]. ∎
@@ 3. Whitney Stratifications from Polar Varieties
The polar approach has two key advantages: first, one can perform all computations in the ring $\mathbb{C}[x_{1},\dots,x_{n}]$ rather than working with the additional $n$ variables $\xi_{1},\dots,\xi_{n}$. And second, one is no longer required to compute any primary decompositions. Both aspects confer significant practical benefits, and the later point will have implications in the complexity analysis presented in Section 6. On the other hand, the new polar algorithm has the disadvantage of being probabilistic [^6] In the sense that its correctness depends on a choice of random constants being truly random, as will be explained later..
@@ 3.1. Polar Varieties and Condition (B)
$L_{\bullet}=\left(L_{0}\supset L_{1}\supset\cdots\supset L_{d-1}\supset L_{d}\right),$
where each $L_{i}\subset\mathbb{C}^{n}$ is an $(n-i)$-dimensional linear subspace.
@@
For each $i$ in ${\left\{{0,1,\ldots,d}\right\}}$, the codimension $i$ polar variety of $X$ along the flag $L_{\bullet}$ is defined as the closure
$\text{\rm P}_{i}(X;L_{\bullet}):=\overline{{\left\{{p\in X_{\text{reg}}\mid\dim(T_{p}X_{\text{reg}}\cap L_{i})\geq d-i}\right\}}}.$
The fact that each $\text{\rm P}_{i}(X;L_{\bullet})$ is an algebraic variety will be established in the next subsection, where we describe explicit generating equations. For now, let us note that if $X$ is irreducible then $\text{\rm P}_{i}(X,L_{\bullet})$ is irreducible for all $i$, that $\text{\rm P}_{0}(X,L_{\bullet})$ always equals $X$, and additionally for a fixed flag of linear spaces $L_{\bullet}$ we obtain a corresponding polar flag
$\text{\rm P}_{d}(X,L_{\bullet})\subset\cdots\subset\text{\rm P}_{0}(X,L_{\bullet})=X.$ (6)
(see e.g. [17, Remark 3.14]).
@@
Our interest in polar varieties stems from a classical result of Teissier relating Condition (B) to polar multiplicities of points of $X$. In order to state this result, let us recall that the Hilbert-Samuel multiplicity of an irreducible subvariety $V\subset X$ is the (integer) coefficient $m_{V}X>0$ of $[X]$ in the Segre class $s(X,V)$ — see [18, Sec 4.3]. These multiplicities may be computed numerically using the algorithm of [22, Theorem 5.3] and even inferred from local point samples as described in [29]. Here is Teissier’s polar multiplicity criterion [17, Theorem 4.13] for Condition (B).
@@ 3.2. Computing (Equations of) Polar Varieties
where the $c^{(i)}_{j}$ are general constants for $i\in{\left\{{1,\dots,d+2}\right\}}$. Let $K_{i}$ be the ideal generated by all $(n-d+i)\times(n-d+i)$ minors of the matrix $\mathscr{K}_{i}$, and let $J$ be the ideal defined by the $(n-d)\times(n-d)$ minors of the matrix $\mathscr{J}$. Then
$\text{\rm P}_{d-i-1}(X,L_{\bullet})=\mathbf{V}_{\mathbb{C}}((K_{i}+I_{X}):J^{\infty}),$ (7)
@@ 3.3. Polar Varieties and Equi-multiplicity
Consider varieties $Y\subset X\subset\mathbb{C}^{n}$, with $X$ a pure dimensional and $Y$ nonempty, irreducible, and different from $P_{i}(X,L_{\bullet})$ for all $i$. Let $\mu=e_{Y}X$ be the Hilbert-Samuel multiplicity of $Y$ in $X$. For all points $z\in Y_{\rm reg}-\text{\rm P}_{r}(X,L_{\bullet})$, we have $e_{z}X=\mu$ where $r=\dim(X)-\dim(Y)$.
@@
First note that multiplicity is unchanged under projective closure, hence we may replace the setup in the statement of the Lemma with $Y\subset X\subset\mathbb{P}^{n}$. By [22, Sec 2.1.4] we may further suppose that $Y=\mathbf{V}_{\mathbb{C}}(f_{1},\ldots,f_{s})$ where all the $f_{i}$ have the same degree. Let $\lambda_{ij}$ be an $r\times s$ matrix of generic complex numbers, and for each $i\in{\left\{{1,\ldots,r}\right\}}$ define the polynomial $P_{i}=\sum_{j}\lambda_{ij}f_{j}$. Setting $W=\mathbf{V}_{\mathbb{C}}(P_{1},\dots,P_{r})$, Theorem 5.3 of [22] gives
$\deg(Y)\mu=\deg(X)\deg(W)-\deg((X\cap W)-Y).$ (8)
Fix a point $z\in Y_{\text{reg}}$, and let $A_{z}\subset\mathbb{P}^{n}$ be a linear space of codimension $\dim(Y)$ which contains $z$. We now examine how (8) changes when all varieties in sight are intersected with $A_{z}$. For generic choices of $A_{z}$, it is known (by [18, Proposition 8.4] for instance) that $\deg(V\cap A_{z})=\deg(V)$ holds for $V\in{\left\{{W,X,Y}\right\}}$. We remark that a sufficient condition for $A_{z}$ to be generic in the sense required here is for $A_{z}$ to be transverse to $X$ in the punctured neighborhood of $z$.
Now, by [28, Proposition 3.3], for generic $A_{z}$ we also have
$e_{z}X=e_{Y\cap A_{z}}(X\cap A_{z}).$
@@
$e_{z}X-\mu=\frac{\deg((X\cap W)-Y)-\deg((X\cap W\cap A_{z})-(Y\cap A_{z}))}{\deg Y}.$ (9)
Next, we seek to establish that the numerator on the right side vanishes unless $z$ lies in $\text{\rm P}_{r}(X;L_{\bullet})$. To this end, let us recall that the numbers ${\left\{{\lambda_{ij}}\right\}}$ which define $W$ are generic. Therefore we may safely assume that $\dim W=n-r$, and that $\dim(X\cap W)=\dim Y$, with the latter equality holding due to transversality. Therefore – at least when $A_{z}$ is generic – both $X\cap W\cap A_{z}$ and $Y\cap A_{z}$ are finite sets of cardinality $\deg(X\cap W)$ and $\deg Y$ respectively. Returning to (9), we conclude that $e_{z}X=\mu$ holds unless $A_{z}$ fails to intersect $X_{\text{reg}}$ transversely near $z$. But since $A_{z}$ might well equal $L_{\dim Y}$ in the flag $L_{\bullet}$, we conclude that $e_{z}X\neq\mu$ forces $z\in\text{\rm P}_{r}(X;L_{\bullet})$. ∎
@@
Output: $X_{\bullet}$, a Whitney stratification of $X$.
@@ 2 Initialize $X_{\bullet}$ to be a flag of length $d+1$ with each ${\rm List}(X_{i}):=\varnothing$
3 For each irreducible component $Z\subset X$, append $Z$ to ${\rm List}(X_{\dim(Z)})$
4 For $i$ from 00 to $d-1$ do
5         For each $j$ append $P_{j}(X_{d-i},L_{\bullet})$ to ${\rm List}(X_{d-i-j})$
6         Set $Y:=(X_{d-i})_{\rm sing}$
7         For each $Z$ an irreducible component of $Y$
8               append $Z$ to ${\rm List}\big{(}X_{\dim(Z)}\big{)}$
9               $V:=Z\cap\text{\rm P}_{d-i-\dim(Z)}(X_{d-i},L_{\bullet})$
10               For each irreducible component $W\subset V$, append $W$ to ${\rm List}\big{(}X_{\dim(W)}\big{)}$
@@ 11 Return $X_{\bullet}$
The algorithm WhitneyPolar above computes a Whitney stratification of $X$.
@@
This is a direct corollary of Lemma 3.5 and Theorem 3.3. In particular, line 5 ensures that within a stratum, each multiplicity in the list in Theorem 3.3 will either be always zero or always non-zero. Lines 6–10 then ensure that multiplicity in the list in Theorem 3.3 is constant via Lemma 3.5. ∎
@@
Note that since $X$ admits a Whitney stratification (by [51, page 336–338] for instance), there exists a proper closed subvariety $Z\subset Y$ such that Whitney’s Condition (B) holds for the pair $(X_{\rm reg},Y_{\rm reg}-Z)$. The reader is warned that the optimal such $Z$, which produces a minimal Whitney stratification, is liable to be much smaller than the subvariety which will be produced by the WhitneyPolar algorithm; this is both becuase the condition we use to achieve constant multiplicity, Lemma 3.5, is a sufficient but not necessary condition, and further since we work globally (rather than locally), along with out polar varieties we inevitably include various random points arising from the choice of $L_{\bullet}$ which are not in fact necessary. We are aware of two antidotes to this — one may either generate a suboptimal stratification and then coarsen it using the algorithm of [25, Sec 3.3], or alternately, one may replace each polar variety in WhitneyPolar with the intersections of several global polar varieties (of the same codimension), each taken along different generic flags $L_{\bullet}$.

### v4

@@ 14 Return $(X_{\bullet},\mathbf{S}_{\bullet}X)$
The termination of WTBStratFlag follows from noting that the recursive call in Line 9 involves the variety $X_{d}$ of dimension at most $d$, with $d$ strictly smaller than $k=\dim X$. And the fact that $X_{\bullet}$ is a Whitney stratification of $X$ follows from the correctness of the original WhitStrat algorithm [24, Sec 4.4]. Moreover, since $\mathbf{F}_{\bullet}X$ is merged into $\mathbf{S}_{\bullet}X$ via Line 7, verifying that $X_{\bullet}$ is $\mathbf{F}_{\bullet}$-subordinate can be reduced to checking that it is $\mathbf{S}_{\bullet}$-subordinate. Therefore, it suffices to confirm that the output pair $(X_{\bullet},\mathbf{S}_{\bullet}X)$ satisfies the two requirements of Definition 2.6. To this end, consider an $i$-dimensional stratum $M$ of $X_{\bullet}$.
@@ 8      Set $\mathbf{F}_{i}X:=f^{-1}(Y^{\prime}_{i})$
[…]
@@ 11 Return $(X_{\bullet},Y_{\bullet})$
Noting that $X^{\prime}_{\bullet}$ is subordinate to the flag $\mathbf{F}_{\bullet}$ by Line 9 of WhitStratMap, we know that for each stratum $M$ of $X^{\prime}_{\bullet}$ there is a number $i:=i(M)$ satisfying $M\subset(\mathbf{F}_{i}X-\mathbf{F}_{i-1}X)$. Now by Line 8, for any such stratum we have $f(M)\subset(Y^{\prime}_{i}-Y^{\prime}_{i-1})$. By the definition of a Whitney stratum we know that $M$ is connected, and so its image under the continuous map $f$ must also be connected; thus there is a unique stratum $N\subset(Y^{\prime}_{i}-Y^{\prime}_{i-1})$ of $Y^{\prime}_{\bullet}$ satisfying $f(M)\subset N$, as claimed above. It remains to establish that the Jacobian of $f|_{M}:M\to N$ has maximal rank at every point of $x$.
@@
Recall that $\mathbf{S}^{\prime\prime}_{j}X$ contains $\mathbf{R}_{j}\overline{M}$ by the second requirement of 2.6 (and similarly for j1j-1). Therefore, the restricted map $f^{*}:\overline{M}_{\rm reg}\to\mathbb{C}^{m}$ has Jacobian of constant rank $\rho_{j}$ given by
@@
Since $M$ is an open subset of $\overline{M}_{\rm reg}$, for each $x\in M$ we immediately have $\text{rank }\big{(}\text{\bf J}f^{*}(x)\big{)}=\rho_{j}.$ Let $\iota:N\hookrightarrow\mathbb{C}^{m}$ be the inclusion map, and recall that $f|_{M}:M\to N$ denotes the restriction of $f$ to $M$, now viewed as a smooth map to $N$ rather than to $\mathbb{C}^{m}$. Thus, $f^{*}=\iota\circ f|_{M}$, and the following diagram of derivatives commutes by the chain rule:
@@ 4      Add $\overline{f(S)}$ to $Y^{+}_{\geq d}$
[…]
@@ 10               Add $Z$ to $X^{\prime}_{\geq r}$
[…]
@@ 14 Return $(X^{\prime}_{\bullet},Y^{\prime}_{\bullet})$
From Theorem 2.7 and Proposition 2.8, we obtain Proposition 2.9 below, which is analogous to [24, Proposition 6.7]. The novelty here lies in the fact that the result below applies to arbitrary polynomial maps between affine complex varieties, and does not impose any genericity requirements. The proof, however, is essentially identical to that of [24, Proposition 6.7].
@@
In [24, Proposition 6.7] this result is obtained for generic mappings $f:X\to Y$, as defined in [24, Definition 6.3], using the construction of the Refine subroutine along with a result ([24, Proposition 6.6]) analogous to that of Proposition 2.8 in the case of generic maps. The new Refine subroutine above is identical to that in [24, Section 6] other than the new version uses WTBStrat at Line 11, rather than WhitStrat. The argument is then identical to that of [24, Proposition 6.7], except we employ the more general result of Proposition 2.8 in place of [24, Proposition 6.6]. ∎
@@ 3. Whitney Stratifications from Polar Varieties
The polar approach has two key advantages: first, one can perform all computations in the ring $\mathbb{C}[x_{1},\dots,x_{n}]$ rather than working with the additional $n$ variables $\xi_{1},\dots,\xi_{n}$. And second, one is no longer required to compute any primary decompositions. Both aspects confer significant practical benefits, as is illustrated in practical tests in Section 6; the later point would also be expected to have significant implications for a worst case complexity analysis of the algorithm (though this is not carried out here). On the other hand, the new polar algorithm has the disadvantage of being probabilistic [^6] In the sense that its correctness depends on a choice of random constants being truly random, as will be explained later..
@@ 3.1. Polar Varieties and Condition (B)
$L_{\bullet}=\left(L_{1}\supset\cdots\supset L_{d-1}\supset L_{d}\right),$
where each $L_{i}\subset\mathbb{C}^{n}$ is an (ni)(n-i)-dimensional linear subspace.
@@
For each $i$ in $\left\{{1,\ldots,d}\right\}$}, the codimension $i$ polar variety of $X$ along the flag $L_{\bullet}$ is defined as the closure
$\text{\rm P}_{i}(X;L_{\bullet}):=\overline{{\left\{{p\in X_{\text{reg}}\mid\dim(T_{p}X_{\text{reg}}\cap L_{i})\geq d-i+1}\right\}}}.$
We define $\text{\rm P}_{0}(X;L_{\bullet}):=X$.
The fact that each $\text{\rm P}_{i}(X;L_{\bullet})$ is an algebraic variety will be established in the next subsection, where we describe explicit generating equations. For now, let us note that if $X$ is irreducible then $\text{\rm P}_{i}(X,L_{\bullet})$ is irreducible for all $i$, and additionally for a fixed flag of linear spaces $L_{\bullet}$ we obtain a corresponding polar flag
$\text{\rm P}_{d}(X,L_{\bullet})\subset\cdots\subset\text{\rm P}_{0}(X,L_{\bullet})=X,$ (7)
(see e.g. [14, Remark 3.14]).
@@
Our interest in polar varieties stems from a classical result of Teissier relating Condition (B) to polar multiplicities of points of $X$. In order to state this result, let us recall that the Hilbert-Samuel multiplicity of an irreducible subvariety $V\subset X$ is the (integer) coefficient $m_{V}X>0$ of $[X]$ in the Segre class $s(X,V)$ — see [15, Sec 4.3]. These multiplicities may be computed numerically using the algorithm of [19, Theorem 5.3] and even inferred from local point samples as described in [27]. Here is Teissier’s polar multiplicity criterion [14, Theorem 4.13] for Condition (B).
@@ 3.2. Computing (Equations of) Polar Varieties
where the $c^{(i)}_{j}$ are general constants for $i\in{\left\{{0,\dots,d-1}\right\}}$. Let $K_{i}$ be the ideal generated by all $(n-d+i+1)\times(n-d+i+1)$ minors of the matrix $\mathscr{K}_{i}$, and let $J$ be the ideal defined by the $(n-d)\times(n-d)$ minors of the matrix $\mathscr{J}$. Then
$\text{\rm P}_{d-i}(X,L_{\bullet})=\mathbf{V}_{\mathbb{C}}((K_{i}+I_{X}):J^{\infty}),$ (8)
@@ 3.3. Polar Varieties and Equi-multiplicity
Consider varieties $Y\subset X\subset\mathbb{C}^{n}$, with $X$ pure dimensional and $Y$ nonempty, irreducible, and different from $\text{\rm P}_{i}(X,L_{\bullet})$ for all $i$. Let $\mu=m_{Y}X$ be the Hilbert-Samuel multiplicity of $Y$ in $X$, and let $r$ be the codimension $\dim X-\dim Y$. Then for all points $z\in Y_{\rm reg}-\text{\rm P}_{r}(X,L_{\bullet})$, we have $m_{z}X=\mu$.
@@
First note that multiplicity is unchanged under projective closure, hence we may replace the setup in the statement of the Lemma with $Y\subset X\subset\mathbb{P}^{n}$. By [19, Sec 2.1.4] we may further suppose that $Y=\mathbf{V}_{\mathbb{C}}(f_{1},\ldots,f_{s})$ where all the $f_{i}$ have the same degree. Let $\lambda_{ij}$ be an $r\times s$ matrix of generic complex numbers, and for each $i\in{\left\{{1,\ldots,r}\right\}}$ define the polynomial $Q_{i}=\sum_{j}\lambda_{ij}f_{j}$. Setting $W=\mathbf{V}_{\mathbb{C}}(Q_{1},\dots,Q_{r})$, Theorem 5.3 of [19] gives
$\deg(Y)\mu=\deg(X)\deg(W)-\deg\left(\overline{(X\cap W)-Y}\right).$ (9)
Fix a point $z\in Y_{\text{reg}}$, and let $A_{z}\subset\mathbb{P}^{n}$ be a linear space of codimension $\dim(Y)$ which contains $z$. We now examine how (9) changes when all varieties in sight are intersected with $A_{z}$. For generic choices of $A_{z}$, it is known (by [15, Proposition 8.4] for instance) that $\deg(V\cap A_{z})=\deg(V)$ holds for $V\in{\left\{{W,X,Y}\right\}}$. We remark that a sufficient condition for $A_{z}$ to be generic in the sense required here is for $A_{z}$ to be transverse to $X$ in the punctured neighborhood of $z$.
Now, by [26, Proposition 3.3], for generic $A_{z}$ we also have
$m_{z}X=m_{Y\cap A_{z}}(X\cap A_{z}).$
@@
$m_{z}X-\mu=\frac{\deg\left(\overline{(X\cap W)-Y}\right)-\deg\left(\overline{(X\cap W\cap A_{z})-(Y\cap A_{z})}\right)}{\deg Y}.$ (10)
Next, we seek to establish that the numerator on the right side vanishes unless $z$ lies in $\text{\rm P}_{r}(X;L_{\bullet})$. To this end, let us recall that the numbers $\left\{{\lambda_{ij}}\right\}$} which define $W$ are generic. Therefore we may safely assume that $\dim W=n-r$, and that $\dim(X\cap W)=\dim Y$, with the latter equality holding due to transversality. Therefore – at least when $A_{z}$ is generic – both $X\cap W\cap A_{z}$ and $Y\cap A_{z}$ are finite sets of cardinality $\deg(X\cap W)$ and $\deg Y$ respectively. Returning to (10), we conclude that $m_{z}X=\mu$ holds unless $A_{z}$ fails to intersect $X_{\text{reg}}$ transversely near $z$. But since $A_{z}$ might well equal $L_{\dim Y}$ in the flag $L_{\bullet}$, we conclude that $m_{z}X\neq\mu$ forces $z\in\text{\rm P}_{r}(X;L_{\bullet})$. ∎
@@
Output: $X_{\bullet}$, the flag in (1) corresponding to the Whitney stratification of $X$.
@@ 2 Initialize $X_{\bullet}$ to be a flag of length $d+1$ with each ${\rm List}(X_{i}):=\varnothing$
3 For each equidimensional component $Z\subset X$, append $Z$ to ${\rm List}(X_{\dim(Z)})$
4 For $i$ from 0 to d1d-1 do
5      For $j$ from $1$ to did-i append $\text{\rm P}_{j}(X_{d-i},L_{\bullet})$ to ${\rm List}(X_{d-i-j})$
6      Compute $Y:=(\bigcup_{j=0}^{i}X_{d-j})_{\rm sing}$
7      For each equidimensional component of $Z\subset Y$, append $Z$ to ${\rm List}\big{(}X_{\dim(Z)}\big{)}$
8      For $j$ from 0 to i1i-1
9         $V:=X_{d-i}\cap\text{\rm P}_{i-j}(X_{d-j},L_{\bullet})$
10         For each equidimensional component $W\subset V$, append $W$ to ${\rm List}\big{(}X_{\dim(W)}\big{)}$
@@ 11 Return $X_{\bullet}$
Remark 3.6.
A recent discussion of algorithms and software to perform equidimensional decomposition (along with a new algorithm for this task) can be found, for example, in [11]. Note that alternatively, depending on the implementation context, one may choose to take irreducible components above everywhere (i.e. in Lines 3, 7, and 10) rather than equidimensional ones.
Remark 3.7.
The WhitneyPolar algorithm presents the flag $X_{\bullet}$ as a union of parts, i.e., a list of components that have the same dimension but are not necessarily irreducible. One may wish to pursue two further goals:
  1. (1)
    design an algorithm to compute the actual connected components in the Whitney stratification, or
  2. (2)
    produce a refinement of the Whitney stratification which ensures that each component is connected.
These tasks depend in practice on whether one works over the real or complex numbers. The latter can be achieved (over $\mathbb{C}$) by decomposing every component in the output into irreducible components.
@@
The algorithm WhitneyPolar is correct.
@@
The first three lines of the algorithm merely populate the flag $X_{\bullet}$ with the equidimensional components of $X$. To prove the correctness of the algorithm, it suffices to show that during the $i$-th iteration of the For loop on Line 4, each (di)(d-i)-dimensional part of the stratification, $\widetilde{X}\in{\rm List}(X_{d-i})$ has been correctly processed. Specifically, we show that at step $i$ the algorithm generates a (finite) list of lower-dimensional subvarieties $Y\subset\widetilde{X}$ such that the union of the generated varieties $Y$ contains all points $y$ which are either not smooth in $X$, or for which the following criterion holds: for every j>dij>d-i and equidimensional component $Z\in{\rm List}(X_{j})$ that contains $\widetilde{X}$, the multiplicity sequence $m_{\bullet}({Z},y)$ differs from the expected sequence $m_{\bullet}({Z},x)$ for generic $x\in\widetilde{X}$.
Line 5 records all points on all polar subvarieties of $\widetilde{X}$; in regards to Lemma˜3.5, this deals separately with the case where $Y$ in the Lemma is a polar subvariety (by ensuring all of these are already removed). Lines 6-7 handle singular points of $\widetilde{X}$. Finally, Lines 8-10 excise the points that are not singular and do not lie on a polar subvariety, but have to be removed to guarantee the conclusion of Lemma˜3.5 (note that we identify a variety containing all points where Lemma˜3.5 can fail for some component of our pure dimensional variety $X_{j}$ all at once, rather than one irreducible component at a time). ∎
@@
The reader is warned that the stratifications produced by WhitneyPolar may not be minimal. There are two reasons for this lack of optimality — first, the condition (from Lemma 3.5) which we use to ensure equi-multiplicity is sufficient but not necessary condition. And second, since our polar varieties are defined globally (rather than locally at a point), they inevitably include various random points arising from the choice of $L_{\bullet}$ which are not in fact necessary. We are aware of two antidotes to this — one may either generate a suboptimal stratification and then coarsen it using the algorithm of [23, Sec 3.3], or alternately, one may replace each polar variety in WhitneyPolar with the intersections of several global polar varieties (of the same codimension), each taken along different generic flags $L_{\bullet}$.